# Strong-coupling quantum thermodynamics using a superconducting flux qubit

### v2

@@ V Conclusion and Outlook
A 675 $\mu$m thick highly resistive Si wafer is used as a substrate. On top of it, a 30 nm thick dielectric film of Al2O3 is deposited using atomic layer deposition (ALD), followed by a 200 nm thick deposition of niobium metal using DC magnetron sputtering. The resonator mask is patterned over a uniform positive e-beam resist using e-beam lithography. After development of patterns, the wafer is baked for a few minutes to strengthen the remaining resist and secure the main structures, avoiding their over-etching. Reactive-ion-etching (RIE) is performed using $SF_{6}$ chemistry to etch the exposed Nb metal. In the last step, the remaining resist is removed using $O_{2}$ descum dry resist stripping.
Figure 8: Thermal voltage to temperature conversion. (a). Typical thermal voltage curve of the SINIS thermometers when biased with varying heating voltages at different cryostat temperatures. (b). Reservoir 1 thermal voltage ($V_{1,\text{Heater}}$), plotted against various cryostat temperatures, at $V_{1,\text{Heater}}=0$, indicated by the dashed line in panel (a). (c-d). The SINIS thermal voltage curve at a cryostat temperature of 100 mK indicates that a voltage of $\approx$ 210 $\mu$V corresponds to a temperature of 250 mK. Black solid arrows are used for better visibility.
For patterning the qubit, a bi-layer resist stack of P(MMA-MMA) copolymer and PMMA is spin-coated, followed by an e-beam exposure. The bi-layer stack is used to create an undercut useful for multi-angle evaporation. We use a solution of Methyl-Isobutyl-Ketone:Isopropanol alcohol (MIBK:IPA volume ratio, 1:3) to develop the PMMA top layer. The undercut profile is provided by a second development step using a Methyl-glycol-Methanol solution with a volume ratio of 1:2. A high-vacuum (HV) e-beam evaporator with an in-situ plasma etching module is used for evaporation. Before evaporation, the sample is exposed to an argon beam for about 40 seconds to clean possible deposited native oxide, enabling a clean contact at Nb-Al interfaces. Using the Dolan bridge shadow evaporation technique [53], a 30 nm thick Al is deposited at $+18^{\circ}$. After the first layer deposition, the devices are oxidized using an in-situ oxidation module (4.3 mbar, for 4 minutes), followed by another Al layer deposition at a second angle ($-18^{\circ}$). Finally, the devices are submerged into a hot $+52^{\circ}$ acetone bath for about an hour to perform the metal lift-off.
For fabricating heat reservoirs and thermometers, which is the next step, we use the three-angle evaporation method. After spin-coating another bi-layer resist stack of P(MMA-MMA) copolymer and PMMA, e-beam exposure, development, and native oxide milling, a 20 nm thick Al layer is deposited at $+39^{\circ}$, forming the superconducting (S) part of the NIS structures. The deposited layer is oxidized under 2.5 mbar chamber pressure for about 5 minutes, forming an insulating tunnel oxide barrier (I). Then a 2 nm thick Al layer is deposited at $-6.5^{\circ}$, followed by the deposition of a 50 nm thick normal metal Cu (N) at the same tilt angle. The 2 nm Al layer makes the adhesion of the Cu layer stronger to the substrate. In the last evaporation step, a 90 nm thick layer of Al is deposited at $+20^{\circ}$. This layer enables a clean contact between the NIS structures and the superconducting resonators.
@@ 1 Thermal Measurement setup
The sample is glued and wire-bonded ultrasonically to a brass sample stage. Encapsulated within two brass caps acting as radiation shields, the sample holder is then plugged to the mixing chamber (MXC) of a custom-made plastic wet dilution refrigerator with base temperature $\approx 70~{}$mK. For read-out purpose, we use conventional thermocoax-filtered cryogenic lines that connect the sample stage from MXC to the room-temperature break-out box. For low-impedance loads, the effective signal bandwidth for the cryogenic filtered lines are between 0-10 kHz. For measurements, programmable voltage sources and Keysight digital multi-meters were employed. For amplification of output voltages we use a room temperature Femto DLPA-100-F-D voltage amplifier. Agilent 33522B waveform generator is used as the bais voltage source. For reading the heat current modulation as a function of applied magnetic field we use SRS SR830 DSP lock-in amplifier which is synchronized to the square-wave modulation of the voltage bias from the Agilent 33522B waveform generator. We perform measurements at several MXC temperatures. The MXC temperatures are controlled by applying voltage to a resistor which is connected to the MXC. For accurate analysis, the calibration of the integrated thermometers (realized as SINIS configuration) were performed at various cryostat temperatures, ranging from 808080~{}80mK up to 360360360~{}360mK. This is done by measuring the voltage drop across the SINIS configuration (set at a current bias of $I_{th}=50~{}$pA), for each cryostat temperature, with no heater bias voltage applied ($V_{1,Heater}=0$V). After setting the MXC to the next higher desired temperature, we wait for over 20 minutes before performing any measurements. This ensures that the device is effectively thermalized to the MXC, allowing us to assume that the phonon temperature is in equilibrium with the MXC temperature. The MXC temperature is monitored using a ruthenium oxide thermometer, calibrated against a Coulomb blockade thermometer. For measurements involving the modulation of the voltage drop across the reservoir thermometers (heat current) as a function of the applied magnetic field, the procedure is as follows: Heat-current measurements as a function of magnetic flux are initiated by stabilizing at the initial flux point for 5 seconds. After this, at each subsequent flux point the system is allowed to stabilize for 2.1 seconds before data acquisition. As previously described, this process is carried out at various values of $T_{0}$, with a waiting time of over 20 minutes after the cryostat temperature is increased to ensure proper thermalization.
@@ 2 Spectroscopy measurement setup
The spectroscopy device, which includes a signal feed line, read-out resonator, and a qubit galvanically coupled to both resonators (as shown in Fig. 5), is mounted on a gold-plated copper sample stage and placed into a dilution refrigerator with a base temperature of $\approx 15~{}$mK. Similar to reported measurement scheme in Ref. [29], the sample holder is equipped with an external superconducting coil for the application of a dc-magnetic field. Suppression of black-body radiation is achieved by employing a series of impedance-matched cryogenic attenuators placed at different temperature stages inside the cryostat. Once attenuated, the input signal passes through the device and is directed through two isolators situated in the cryostat mixing chamber. Subsequently, it undergoes amplification by a 424242~{}42dB low-noise high-electron-mobility transistor (HEMT) amplifier at the 4-K stage, followed by further amplification of 525252~{}52dB using two additional amplifiers positioned at room temperature. For single-tone spectroscopy, we utilize a commercially available vector network analyzer (VNA). For two-tone spectroscopy, we use a commercial microwave signal generator. The frequency range of interest, 4-8 GHz, is supported by the setup. The measurement setup and procedures for single-tone and two-tone spectroscopy are similar to those outlined in Ref. [29].
@@ 1 Thermal Measurements
A custom made, plastic wet dilution refrigerator is used for thermal conductance measurements. These measurements are conducted over a mixing chamber (MXC) temperature range of $80$ to 360360360~{}360mK, which is controlled by applying voltage to a resistor located at MXC of the cryostat. The device achieves effective thermalization with the MXC, allowing us to assume that the phonon temperature, $T_{\rm 0}$, is in equilibrium with that of the MXC. This temperature is accurately measured using a ruthenium oxide thermometer, which has been calibrated with a Coulomb blockade thermometer for precision. The detailed measurement setup configuration is provided in Appendix B. Reservoir 1 is locally heated using an NIS junction which is biased by an external DC voltage source $V_{\rm{1,\,Heater}}$. The Reservoir 1 and Reservoir 2 thermometers are biased with a small current of $I_{th}=50~{}$pA, and the voltage drop across the thermometers ($V_{\rm{1,\,Th}}$ and $V_{\rm{2,\,Th}}$, respectively), is recorded at several cryostat temperatures $T_{0}$. To determine the temperatures of the Reservoir 1 ($T_{\text{1}}$) and Reservoir 2 ($T_{\text{2}}$), we measure the thermal voltage drop across both thermometers (set at a current bias of $I_{th}=50~{}$pA), for each cryostat temperature ($T_{\text{0}}$), with no heating bias voltage applied ($V_{\rm{1,\,Heater}}=0$). The recorded thermal voltage is converted to its respective temperature using the calibration method described in Appendix B. Using the calibrated data, the absorbed power by Reservoir 2, $P_{1\to 2}$ ,is then calculated using expression, $P_{\rm ep,2}=\Sigma\mathcal{V}\,(T_{\rm 2}^{5}-T_{\rm 0}^{5})$. The magnetic flux controlling the energy of the qubit is provided by a external superconducting coil located at 4 K temperature. To observe the modulation in the voltage drop of thermometers, $V_{\rm{2,\,Th}}$, at Reservoir 2, we heat the Reservoir 1 ($V_{\rm{1,\,Heater}}\neq 0$) simultaneously sweeping the magnetic field, and record the change in $V_{\rm{2,\,Th}}$. After calibrating the data, we arrive at the figure presented in Fig. 3.
@@ 2 Thermal voltage to Power conversion: Mechanism
Figure 8 (a) presents the measurement of the thermometer voltage, $V_{\rm{1,\,Th}}$, at a fixed applied current of $I_{\rm{1,Th}}=50$ pA for various bath temperatures, $T_{0}=80-360$ mK, with a temperature interval of 20 mK between each measurement, from top to bottom, against the voltage, $V_{\rm{1,\,Heater}}$, across the NIS heater. The equilibrium values of $V_{\rm{1,\,Th}}$ at $V_{\rm{1,\,Heater}}=0$, represented by symbols in The equilibrium values of $V_{\rm{1,\,Th}}$ at $V_{\rm{1,\,Heater}}=0$, represented by symbols in Fig. 8 (b) (with the position indicated by the black dashed line in panel (a)), are plotted as a function of bath temperature, $T_{0}$. These data points $V_{\rm{1,\,Th}}(V_{\rm{1,\,Heater}}=0)$ serve as the calibration for the corresponding reservoir, here the Reservoir 1 temperature. The solid black line represents a fit to the measured data. At $80$ mK $\leq T_{1}<135$ mK, the fitting function is $T_{1}=a+b*\lg(c-V_{\rm{1,\,Th}})$ with $a$, $b$ and $c$ as constant [54], and at 135 mK $\leq T_{1}<370$ mK we used standard third-order polynomial (nearly linear) function. Panels 8 (c) and 8 (d) display how the Reservoir 1 temperature $T_{1}$ is extracted. With the same procedure we can obtain the Reservoir 2 temperature, $T_{2}$. The measured temperatures are then converted to power using Eq. (1) described in the main text to obtain the power transmitted to the Reservoir 2.
@@ 1 Flux qubit hybridized with two resonators coupled to two reservoirs
[…]
where ^$\hat{\mathcal{H}}_{\rm q}$ is the Hamiltonian of the qubit, ^$\hat{\mathcal{H}}_{i}$ that of each resonator, ^$\hat{\mathcal{H}}_{{\rm c}i}$ the coupling between the resonator $i$ and the qubit, and ^$\hat{\mathcal{H}}_{\rm 12}$ the coupling between the two resonators. We write Eq. (7) in the form [48]
[…]
Here, $\hat{b}^{\dagger}$ and $\hat{b}$ represent the creation and annihilation operators for the qubit, and similarly, $\hat{a}^{\dagger}_{i}$ and $\hat{a}_{i}$ are for resonator $i$, $g_{i}$ denotes the coupling constant between the qubit and resonator $i$, and $\gamma$ between the two resonators. ^$\hat{\mathcal{H}}_{\rm q}$ can be written as
^$\hat{\mathcal{H}}_{\rm q}=hf_{\rm q}(\Phi)|e\rangle\langle e|.$ (9)
@@
[…]
@@
^$\frac{\partial\hat{\mathcal{H}}_{\rm qci}}{\partial\Phi}=-h\,g_{i}(\hat{a}_{i}+\hat{a}_{i}^{\dagger})\bigg{(}\frac{\partial|e\rangle}{\partial\Phi}\langle g|+|e\rangle\frac{\partial\langle g|}{\partial\Phi}+\frac{\partial|g\rangle}{\partial\Phi}\langle e|+|g\rangle\frac{\partial\langle e|}{\partial\Phi}\bigg{)},$ (38)
^$\frac{\partial\hat{\mathcal{H}}_{\rm qci}}{\partial\Phi}=\frac{g_{i}}{\Phi_{0}}\frac{I_{p}}{f_{\rm q}^{(0)}}\,\bigg{(}\frac{f_{\rm q}^{(0)}}{f_{\rm q}}\bigg{)}^{2}(\hat{a}_{i}+\hat{a}_{i}^{\dagger})(2\hat{b}^{\dagger}\hat{b}-1),$ (39)
@@ 2 Coupling of the noise to the flux qubit
where $\Gamma_{i\rightarrow j}=\Gamma_{i\rightarrow j}^{r=1}+\Gamma_{i\rightarrow j}^{r=2}$, where the superscript $r$ refers to the both Reservoir 1 and Reservoir 2. For steady state we put $\dot{\rho}_{i}=0$, $(i=0...3)$, and $\sum_{i=0}^{3}\rho_{i}=1$, which yields the populations of the levels ${\rho}_{i}$ as
Figure 12: Observations in theory and experimental data. (a). The experimental results (left) and the theoretical prediction (right) for the power transmitted in the setup. The parameters in the calculation are as follows: $R=6\,\Omega$, $L=0.8\,$nH, $I_{p}=30\,$nA, $f_{\rm q}^{(0)}=2\,$GHz, and $f_{\rm r}=7\,$GHz.

### v3

@@ V Conclusion and Outlook
[…]
For patterning the qubit, a bi-layer resist stack of P(MMA-MMA) copolymer and PMMA is spin-coated, followed by an e-beam exposure. The bi-layer stack is used to create an undercut useful for multi-angle evaporation. We use a solution of Methyl-Isobutyl-Ketone:Isopropanol alcohol (MIBK:IPA volume ratio, 1:3) to develop the PMMA top layer. The undercut profile is provided by a second development step using a Methyl-glycol-Methanol solution with a volume ratio of 1:2. A high-vacuum (HV) e-beam evaporator with an in-situ plasma etching module is used for evaporation. Before evaporation, the sample is exposed to an argon beam for about 40 seconds to clean possible deposited native oxide, enabling a clean contact at Nb-Al interfaces. Using the Dolan bridge shadow evaporation technique [53], a 30 nm thick Al is deposited at $18^{\circ}$. After the first layer deposition, the devices are oxidized using an in-situ oxidation module (4.3 mbar, for 4 minutes), followed by another Al layer deposition at a second angle ($18^{\circ}$). Finally, the devices are submerged into a hot $52^{\circ}$C acetone bath for about an hour to perform the metal lift-off.
For fabricating heat reservoirs and thermometers, which is the next step, we use the three-angle evaporation method. After spin-coating another bi-layer resist stack of P(MMA-MMA) copolymer and PMMA, e-beam exposure, development, and native oxide milling, a 20 nm thick Al layer is deposited at $39^{\circ}$, forming the superconducting (S) part of the NIS structures. The deposited layer is oxidized under 2.5 mbar chamber pressure for about 5 minutes, forming an insulating tunnel oxide barrier (I). Then a 2 nm thick Al layer is deposited at $6.5^{\circ}$, followed by the deposition of a 50 nm thick normal metal Cu (N) at the same tilt angle. The 2 nm Al layer makes the adhesion of the Cu layer stronger to the substrate. In the last evaporation step, a 90 nm thick layer of Al is deposited at $20^{\circ}$. This layer enables a clean contact between the NIS structures and the superconducting resonators.
@@ 1 Thermal Measurement setup
The sample is glued and wire-bonded ultrasonically to a brass sample stage. Encapsulated within two brass caps acting as radiation shields, the sample holder is then plugged to the mixing chamber (MXC) of a custom-made plastic wet dilution refrigerator with base temperature $\approx 70$~mK. For read-out purpose, we use conventional thermocoax-filtered cryogenic lines that connect the sample stage from MXC to the room-temperature break-out box. For low-impedance loads, the effective signal bandwidth for the cryogenic filtered lines are between 0-10 kHz. For measurements, programmable voltage sources and Keysight digital multi-meters were employed. For amplification of output voltages we use a room temperature Femto DLPA-100-F-D voltage amplifier. Agilent 33522B waveform generator is used as the bais voltage source. For reading the heat current modulation as a function of applied magnetic field we use SRS SR830 DSP lock-in amplifier which is synchronized to the square-wave modulation of the voltage bias from the Agilent 33522B waveform generator. We perform measurements at several MXC temperatures. The MXC temperatures are controlled by applying voltage to a resistor which is connected to the MXC. For accurate analysis, the calibration of the integrated thermometers (realized as SINIS configuration) were performed at various cryostat temperatures, ranging from 8080~mK up to 360360~mK. This is done by measuring the voltage drop across the SINIS configuration (set at a current bias of $I_{th}=50$~pA), for each cryostat temperature, with no heater bias voltage applied ($V_{1,Heater}=0$V). After setting the MXC to the next higher desired temperature, we wait for over 20 minutes before performing any measurements. This ensures that the device is effectively thermalized to the MXC, allowing us to assume that the phonon temperature is in equilibrium with the MXC temperature. The MXC temperature is monitored using a ruthenium oxide thermometer, calibrated against a Coulomb blockade thermometer. For measurements involving the modulation of the voltage drop across the reservoir thermometers (heat current) as a function of the applied magnetic field, the procedure is as follows: Heat-current measurements as a function of magnetic flux are initiated by stabilizing at the initial flux point for 5 seconds. After this, at each subsequent flux point the system is allowed to stabilize for 2.1 seconds before data acquisition. As previously described, this process is carried out at various values of $T_{0}$, with a waiting time of over 20 minutes after the cryostat temperature is increased to ensure proper thermalization.
@@ 2 Spectroscopy measurement setup
The spectroscopy device, which includes a signal feed line, read-out resonator, and a qubit galvanically coupled to both resonators (as shown in Fig. 5), is mounted on a gold-plated copper sample stage and placed into a dilution refrigerator with a base temperature of $\approx 15$~mK. Similar to reported measurement scheme in Ref. [29], the sample holder is equipped with an external superconducting coil for the application of a dc-magnetic field. Suppression of black-body radiation is achieved by employing a series of impedance-matched cryogenic attenuators placed at different temperature stages inside the cryostat. Once attenuated, the input signal passes through the device and is directed through two isolators situated in the cryostat mixing chamber. Subsequently, it undergoes amplification by a 4242~dB low-noise high-electron-mobility transistor (HEMT) amplifier at the 4-K stage, followed by further amplification of 5252~dB using two additional amplifiers positioned at room temperature. For single-tone spectroscopy, we utilize a commercially available vector network analyzer (VNA). For two-tone spectroscopy, we use a commercial microwave signal generator. The frequency range of interest, 4-8 GHz, is supported by the setup. The measurement setup and procedures for single-tone and two-tone spectroscopy are similar to those outlined in Ref. [29].
@@ 1 Thermal Measurements
A custom made, plastic wet dilution refrigerator is used for thermal conductance measurements. These measurements are conducted over a mixing chamber (MXC) temperature range of $80$ to 360360~mK, which is controlled by applying voltage to a resistor located at MXC of the cryostat. The device achieves effective thermalization with the MXC, allowing us to assume that the phonon temperature, $T_{\rm 0}$, is in equilibrium with that of the MXC. This temperature is accurately measured using a ruthenium oxide thermometer, which has been calibrated with a Coulomb blockade thermometer for precision. The detailed measurement setup configuration is provided in Appendix B. Reservoir 1 is locally heated using an NIS junction which is biased by an external DC voltage source $V_{\rm{1,\,Heater}}$. The Reservoir 1 and Reservoir 2 thermometers are biased with a small current of $I_{th}=50$~pA, and the voltage drop across the thermometers ($V_{\rm{1,\,Th}}$ and $V_{\rm{2,\,Th}}$, respectively), is recorded at several cryostat temperatures $T_{0}$. To determine the temperatures of the Reservoir 1 ($T_{\text{1}}$) and Reservoir 2 ($T_{\text{2}}$), we measure the thermal voltage drop across both thermometers (set at a current bias of $I_{th}=50$~pA), for each cryostat temperature ($T_{\text{0}}$), with no heating bias voltage applied ($V_{\rm{1,\,Heater}}=0$). The recorded thermal voltage is converted to its respective temperature using the calibration method described in Appendix B. Using the calibrated data, the absorbed power by Reservoir 2, $P_{1\to 2}$ ,is then calculated using expression, $P_{\rm ep,2}=\Sigma\mathcal{V}\,(T_{\rm 2}^{5}-T_{\rm 0}^{5})$. The magnetic flux controlling the energy of the qubit is provided by a external superconducting coil located at 4 K temperature. To observe the modulation in the voltage drop of thermometers, $V_{\rm{2,\,Th}}$, at Reservoir 2, we heat the Reservoir 1 ($V_{\rm{1,\,Heater}}\neq 0$) simultaneously sweeping the magnetic field, and record the change in $V_{\rm{2,\,Th}}$. After calibrating the data, we arrive at the figure presented in Fig. 3.
@@ 2 Thermal voltage to Power conversion: Mechanism
Figure 8 (a) presents the measurement of the thermometer voltage, $V_{\rm{1,\,Th}}$, at a fixed applied current of $I_{\rm{1,Th}}=50$ pA for various bath temperatures, $T_{0}=80-360$ mK, with a temperature interval of 20 mK between each measurement, from top to bottom, against the voltage, $V_{\rm{1,\,Heater}}$, across the NIS heater. The equilibrium values of $V_{\rm{1,\,Th}}$ at $V_{\rm{1,\,Heater}}=0$, represented by symbols in Fig. 8 (b) (with the position indicated by the black dashed line in panel (a)), are plotted as a function of bath temperature, $T_{0}$. These data points $V_{\rm{1,\,Th}}(V_{\rm{1,\,Heater}}=0)$ serve as the calibration for the corresponding reservoir, here the Reservoir 1 temperature. The solid black line represents a fit to the measured data. At $80$ mK $\leq T_{1}<135$ mK, the fitting function is $T_{1}=a+b*\lg(c-V_{\rm{1,\,Th}})$ with $a$, $b$ and $c$ as constant [54], and at 135 mK $\leq T_{1}<370$ mK we used standard third-order polynomial (nearly linear) function. Panels 8 (c) and 8 (d) display how the Reservoir 1 temperature $T_{1}$ is extracted. With the same procedure we can obtain the Reservoir 2 temperature, $T_{2}$. The measured temperatures are then converted to power using Eq. (1) described in the main text to obtain the power transmitted to the Reservoir 2.
@@ 1 Flux qubit hybridized with two resonators coupled to two reservoirs
$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm q}+\sum_{i=1,\,2}\hat{\mathcal{H}}_{i}+\sum_{i=1,\,2}\hat{\mathcal{H}}_{{\rm c}i}+\hat{\mathcal{H}}_{\rm 12},$ (7)
where $\hat{\mathcal{H}}_{\rm q}$ is the Hamiltonian of the qubit, $\hat{\mathcal{H}}_{i}$ that of each resonator, $\hat{\mathcal{H}}_{{\rm c}i}$ the coupling between the resonator $i$ and the qubit, and $\hat{\mathcal{H}}_{\rm 12}$ the coupling between the two resonators. We write Eq. (7) in the form [48]
$\hat{\mathcal{H}}=hf_{\rm q}(\Phi)\hat{b}^{\dagger}\hat{b}+\sum_{i=1,\,2}hf_{i}\hat{a}_{i}^{\dagger}\hat{a}_{i}-\sum_{i=1,\,2}hg_{j}(\hat{a}_{j}+\hat{a}_{j}^{\dagger})(\hat{b}+\hat{b}^{\dagger})-h\gamma(\hat{a}_{1}+\hat{a}_{1}^{\dagger})(\hat{a}_{2}+\hat{a}_{2}^{\dagger}).$ (8)
Here, $\hat{b}^{\dagger}$ and $\hat{b}$ represent the creation and annihilation operators for the qubit, and similarly, $\hat{a}^{\dagger}_{i}$ and $\hat{a}_{i}$ are for resonator $i$, $g_{i}$ denotes the coupling constant between the qubit and resonator $i$, and $\gamma$ between the two resonators. $\hat{\mathcal{H}}_{\rm q}$ can be written as
$\hat{\mathcal{H}}_{\rm q}=hf_{\rm q}(\Phi)|e\rangle\langle e|.$ (9)
@@
where $\hbar\omega_{ij}$ is the energy separation between states $i$ and $j$. In the Hamiltonian of Eq. (7), there are two contributions, $\hat{\mathcal{H}}_{\rm q}$ and $\sum_{i=1,\,2}\hat{\mathcal{H}}_{\rm qci}$, that depend on flux $\Phi$. The derivative of each of them with respect to flux is presented below. For $\hat{\mathcal{H}_{\rm q}}$ we have
@@
$\frac{\partial\hat{\mathcal{H}}_{\rm qci}}{\partial\Phi}=-h\,g_{i}(\hat{a}_{i}+\hat{a}_{i}^{\dagger})\bigg(\frac{\partial|e\rangle}{\partial\Phi}\langle g|+|e\rangle\frac{\partial\langle g|}{\partial\Phi}+\frac{\partial|g\rangle}{\partial\Phi}\langle e|+|g\rangle\frac{\partial\langle e|}{\partial\Phi}\bigg),$ (38)
$\frac{\partial\hat{\mathcal{H}}_{\rm qci}}{\partial\Phi}=\frac{g_{i}}{\Phi_{0}}\frac{I_{p}}{f_{\rm q}^{(0)}}\,\bigg(\frac{f_{\rm q}^{(0)}}{f_{\rm q}}\bigg)^{2}(\hat{a}_{i}+\hat{a}_{i}^{\dagger})(2\hat{b}^{\dagger}\hat{b}-1),$ (39)
@@ 2 Coupling of the noise to the flux qubit
where $\Gamma_{i\rightarrow j}=\Gamma_{i\rightarrow j}^{r=1}+\Gamma_{i\rightarrow j}^{r=2}$, where the superscript $r$ refers to the both Reservoir 1 and Reservoir 2. For steady state we put $\dot{\rho}_{i}=0$, (i=03)(i=0...3), and $\sum_{i=0}^{3}\rho_{i}=1$, which yields the populations of the levels ${\rho}_{i}$ as
Figure 12: Observations in theory and experimental data. (a). The experimental results (left) and the theoretical prediction (right) for the power transmitted in the setup. The parameters in the calculation are as follows: $R=6\,\Omega$, L=0.8L=0.8\,nH, $I_{p}=30\,$nA, $f_{\rm q}^{(0)}=2\,$GHz, and $f_{\rm r}=7\,$GHz.
@@
3 Flux qubit coupled to a bare resistor
We can obtain the power nearly quantitatively at the symmetry point $\Phi=\Phi_{0}/2$, i.e. “ the central peak” by simply assuming that a bare resistor (reservoir) is coupled inductively to the qubit.
Figure 13: Thermal noise of a resistor to a flux qubit via inductive coupling.
The equivalent circuit representing a resistor with resistance $R$ coupled via mutual inductance $M$ to a flux qubit is shown in Fig. 13. Based on the Kirchhoff’s circuit law, the current $I$ reads
$I=\frac{R}{R+\imath\omega L}I_{n}$ (52)
The noise spectral density of (52) is then given by
$S_{I}(\omega)=\frac{R^{2}}{R^{2}+\omega^{2}L^{2}}S_{I_{n}}(\omega).$ (53)
Since for the bare resistor with temperature $T=(k_{B}\beta)^{-1}$, $S_{I_{n}}(\omega)=\frac{2}{R}\frac{\hbar\omega}{1-e^{-\beta\hbar\omega}}$, we have
$S_{I}(\omega)=\frac{2R}{R^{2}+\omega^{2}L^{2}}\frac{\hbar\omega}{1-e^{-\beta\hbar\omega}}.$ (54)
The flux into the qubit is $\Phi=M\,I$. Using Eq. (54), the noise sepectral density of flux, $S_{\Phi}=M^{2}\,S_{I}$, reads
$S_{\Phi}(\omega)=2M^{2}\frac{R}{R^{2}+\omega^{2}L^{2}}\frac{\hbar\omega}{1-e^{-\beta\hbar\omega}}.$ (55)
Based on the golden rule, transition rate up, $\Gamma_{\uparrow}$, between the two levels of the qubit (ground state $|g\rangle$ and excited state $|e\rangle$) due to the presence of the bath (resistor $R$) is given by
$\Gamma_{\uparrow}=\frac{1}{\hbar^{2}}|\langle g|\frac{\partial H}{\partial\Phi}|e\rangle|^{2}S_{\Phi}(-\omega_{01}),$ (56)
[…]
$\Gamma_{\uparrow}=\frac{I_{p}^{2}M^{2}}{\hbar^{2}}\frac{2R}{R^{2}+\omega_{01}^{2}L^{2}}\frac{\hbar\omega_{01}}{1-e^{-\beta\hbar\omega_{01}}}.$ (57)
In order to obtain the power between the two baths at temperatures $T_{1}$ and $T_{2}$, we need the diagonal elements of the density matrix. With no time-dependent driving we then have
$\dot{\rho}_{ee}=\Gamma_{\Sigma,\uparrow}\rho_{gg}-\Gamma_{\Sigma,\downarrow}\rho_{ee},$ (58)
where $\Gamma_{\Sigma,\uparrow}=\Gamma_{1,\uparrow}+\Gamma_{2,\uparrow}$ and $\Gamma_{\Sigma,\downarrow}=\Gamma_{1,\downarrow}+\Gamma_{2,\downarrow}$. By applying steady-state condition $\dot{\rho}_{ee}=0$, the population of the excited state reads
${\rho}_{ee}=\frac{\Gamma_{\Sigma,\uparrow}}{\Gamma_{\Sigma,\downarrow}+\Gamma_{\Sigma,\uparrow}}.$ (59)
The power, $\dot{Q}$, that goes out from the bath (resistor) is given by
$\dot{Q}=\hbar\omega_{01}[\Gamma_{1,\uparrow}{\rho}_{gg}-\Gamma_{1,\downarrow}{\rho}_{ee}],$ (60)
where ${\rho}_{gg}=1-{\rho}_{ee}$ denotes the population of the ground state of the qubit. If we assume a two-bath system with $T_{2}\ll T_{1}$, then this leads to $\Gamma_{2,\uparrow}\simeq 0$, and $\Gamma_{2,\downarrow}\simeq\Gamma_{1,\downarrow}$. In this case power $\dot{Q}$ reads
$\dot{Q}$ $\simeq\frac{1}{2}\hbar\omega_{01}\Gamma_{1,\uparrow},$ (61)
We then have
$\dot{Q}\simeq I_{p}^{2}M^{2}\frac{R}{R^{2}+\omega_{01}^{2}L^{2}}\frac{\omega_{01}^{2}}{e^{\beta_{1}\hbar\omega_{01}}-1},$ (62)
where $\beta_{1}=(k_{B}T_{1})^{-1}$ is the temperature of the hot bath (resistor). For the galvanic coupling here as we have here, $M\simeq L$. We then have
$\dot{Q}\simeq I_{p}^{2}R\frac{(\omega_{01}L)^{2}}{R^{2}+(\omega_{01}L)^{2}}\frac{1}{e^{\beta_{1}\hbar\omega_{01}}-1},$ (63)
We see that for $L\ll R/\omega_{01}$, $\dot{Q}\propto L^{2}$ and it obtains its maximum value
$\dot{Q}_{\rm max}=I_{p}^{2}L\frac{\omega_{01}}{e^{\beta_{1}\hbar\omega_{01}}-1},$ (64)
when $L=R/\omega_{01}$. With the values used in the theoretical model, with the parameter values as stated above, we have from Eq. (63), $\dot{Q}\simeq 7.5$~fW at $T_{1}=300$ mK, which remains in agreement with the experiment.